| 
Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
FXE
SPB
HED
SQS
SXP
Sample Environment
PSPO
XO
TS
Migrated2Zulip
Archived2GitlabPages
Archived2PDF
MigratedByGroup
 | 
| 
SASE1 Status and SASE3 Readiness
 | 
|   Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther  | 
Not logged in | 
  | 
   | 
| 
Message ID: 515
    Entry time: 23 Apr 2018, 15:35
 |  
| Author: | 
Adriano Violante  | 
 | Location: | 
General  | 
 | Severity: | 
Information  | 
 | Subject: | 
Status report after dispatch meeting 13:00  | 
 
  |  
 
 | 
SPB 
	- Timing tool for optical laser installed
 
	- Sample chamber's PLC will be updated next week due to ongoing tests until monday. Pending.
 
	- It is not clear whether the AGIPD firmware has to be updated. To clarify with Jola when she is back. The info is important foe ITDM to plan the DAQ tests. Pending. 
 
	- During the last users run it happened that a valve in XTD9 was closed but the beam did not trip. It is not clear why. Will be clarified with AE. Pending.
 
 
AE 
	- SASE2 EPS Loop running
 
	- Updating PLCs for MPS in the tunnels. The update was stopped last friday due to a problem with the on/off funcionality of some ion pumps. The problem was found and the fixing requires an update of the VAC loops whoch will be done in the next hours/days. In case the loop update is not possible due to specific requirements the problem can be bypassed in Karabo. 
 
 
  
General 
On Wednesday the meeting will be held at 13:00. Starting from next monday it will go back to 9.30  |